# QoE Maximization for Multiple-UAV-Assisted Multi-Access Edge Computing: An Online Joint Optimization Approach

### v1

@@ 5.1 Stage 1: Task Offloading and Resource Allocation
Problem P1 is still a MINLP problem, and the decisions of task offloading and resource allocation are coupled with each other. Considering that the UAVs are service providers in the considered UAV-assisted MEC system, we prioritize resource allocation strategies for the UAVs. Then, based on the resource allocation strategies, we optimize the offloading decisions of UDs.
5.1.1 Resource Allocation
Given an arbitrary task offloading decision profile $\mathcal{A}$ of the UDs, the servers decides resource allocation strategies to minimize problem P1. Define $z_{s,m}=\frac{F_{s,m}}{F_{s}^{\text{max}}}$ and remove irrelevant constant terms, the resource allocation problem can be formulated as
@@ 5.1.1 Resource Allocation
To solve problem P1.1, we first prove that the problem is a convex optimization problem through Lemma 1. Then, based on Lemma 1, we can obtain the optimal resource allocation by Theorem 2.
@@
First, we define $F(\mathcal{Z},\mathcal{W})$ to denote the objective function (32). Taking the second-order partial derivative of $F(\mathcal{S},\mathcal{W})$, we have
$\small\begin{cases}\frac{\partial^{2}F}{\partial(z_{s,m})^{2}}=\frac{2\gamma_{m}^{\text{T}}\eta_{m}D_{m}}{(z_{s,m})^{3}F_{s}^{\mathrm{max}}},\\ \frac{\partial^{2}F}{\partial(w_{s,m})^{2}}=\frac{2(\gamma_{m}^{\text{T}}D_{m}+\gamma_{m}^{\text{E}}P_{m}D_{m})}{(w_{s,m})^{3}r_{s,m}}.\end{cases}$ (33)
Because $\frac{\partial^{2}F}{\partial(z_{s,m})^{2}}\geq 0$ and $\frac{\partial^{2}F}{\partial(w_{s,m})^{2}}\geq 0$, the Hessian matrix of $F(\mathcal{S},\mathcal{W})$ is positive semidefinite. Furthermore, the constraints (32a)-(32d) are linear. Therefore, problem P1.1 is a convex optimization problem.
@@
First, the Lagrange function of Problem P1.1 can be defined as:
$L(\mathcal{Z},\mathcal{W},\mathbf{\alpha},\beta,\lambda,\sigma)=F(\mathcal{Z},\mathcal{W})-\sum_{s\in\mathcal{S}}\sum_{m\in\mathbf{M}_{s}(\mathcal{A})}\alpha_{s,m}z_{s,m}$
$-\sum_{s\in\mathcal{S}}\sum_{m\in\mathbf{M}_{s}(\mathcal{A})}\beta_{s,m}w_{s,m}+\sum_{s\in\mathcal{S}}\lambda_{s}(\sum_{m\in\mathbf{M}_{s}(\mathcal{A})}z_{s,m}-1)$
$+\sum_{s\in\mathcal{S}}\sigma_{s}(\sum_{m\in\mathbf{M}_{s}(\mathcal{A})}w_{s,m}-1),$ (35)
where $\mathbf{\alpha}=\{\alpha_{s,m}\}_{s\in\mathcal{S},m\in\mathbf{M}_{s}(\mathcal{A})}$, $\mathbf{\beta}=\{\beta_{s,m}\}_{s\in\mathcal{S},m\in\mathbf{M}_{s}(\mathcal{A})}$, $\lambda=\{\lambda_{s}\}_{s\in\mathcal{S}}$ and $\sigma=\{\sigma_{s}\}_{s\in\mathcal{S}}$ represent the Lagrange multipliers of constraints (32a)-(32d) and are non-negative, respectively. Then, we can obtain the following Karush–Kuhn–Tucker (KKT) conditions.
Stationarity: $\frac{\partial L(\mathcal{Z}^{*},\mathcal{W}^{*},\mathbf{\alpha}^{*},\beta^{*},\mathbf{\lambda}^{*},\mathbf{\sigma}^{*})}{\partial z_{s,m}}=0$,
$\frac{\partial L(\mathcal{Z}^{*},\mathcal{W}^{*},\mathbf{\alpha}^{*},\beta^{*},\mathbf{\lambda}^{*},\mathbf{\sigma}^{*})}{\partial w_{s,m}}=0$,
$\forall s\in\mathcal{S},m\in\mathbf{M}_{s}(\mathcal{A})$
Primal feasibility: $z_{s,m}^{*}\geq 0,\forall s\in\mathcal{S},\ m\in\mathbf{M}_{s}(\mathcal{A})$
$\sum_{m\in\mathbf{M}_{s}(\mathcal{A})}z_{s,m}^{*}\leq 1,\forall s\in\mathcal{S}$
$w_{s,m}^{*}\geq 0,\forall s\in\mathcal{S},\ m\in\mathbf{M}_{s}(\mathcal{A})$
$\sum_{m\in\mathbf{M}_{s}(\mathcal{A})}w_{s,m}^{*}\leq 1,\forall s\in\mathcal{S}$
Dual feasibility: $\alpha_{s,m}^{*},\beta_{s,m}^{*},\lambda_{s}^{*},\sigma_{s}^{*}\geq 0,\forall s\in\mathcal{S},m\in\mathbf{M}_{s}(\mathcal{A})$
Complementary $-\alpha_{s,m}^{*}z_{s,m}^{*}=0,\forall s\in\mathcal{S},m\in\mathbf{M}_{s}(\mathcal{A})$
slackness: $-\beta_{s,m}^{*}w_{s,m}^{*}=0,\forall s\in\mathcal{S},m\in\mathbf{M}_{s}(\mathcal{A})$
$\lambda_{s}^{*}(\sum_{m\in\mathbf{M}_{s}(\mathcal{A})}z_{s,m}^{*}-1)=0,\forall s\in\mathcal{S}$
$\sigma_{s}^{*}(\sum_{m\in\mathbf{M}_{s}(\mathcal{A})}w_{s,m}^{*}-1)=0,\forall s\in\mathcal{S}$
Based on Lemma 1, i.e., problem P1.1 is a convex optimization problem, the solution that satisfies the KKT condition mentioned above is also the optimal solution of problem P1.1. Solving the above KKT conditions, we can obtain the optimal resource allocation coefficient that is defined in (34).
5.1.2 Task Offloading
@@
Therefore, we can design the utility function of UD $m$ as follows:
@@
However, the task offloading decision for UD $m$ is influenced by not only its individual demand but also the offloading decisions made by other UDs. Given the competitive dynamics inherent in task offloading among UDs, we utilize game theory to address this subproblem.
@@
  • $\mathbb{A}=\mathbf{A}_{1}\times\dots\times\mathbf{A}_{M}$ denotes the strategy space, wherein $\mathbf{A}_{m}=\{0\}\cup\{u\}\cup{\mathcal{N}}$ is the set of offloading strategies for player $m\ (m\in\mathcal{M})$, $a_{m}\in\mathbf{A}_{m}$ denotes the offloading decision of player $m$, and $\mathcal{A}=(a_{1},\dots,a_{M})\in\mathbb{A}$ is the strategy profile.
  • $(U_{m}(\mathcal{A}))_{m\in\mathcal{M}}$ represents the utility of player $m$, which maps each strategy profile $\mathcal{A}$ to a real number.
Every player strives to minimize its utility by selecting an appropriate offloading strategy. Therefore, the MUD-TOG can be described by the following distributed optimization problem mathematically:
@@
(2) The solution to MUD-TOG. For MUD-TOG, the concept of Nash equilibrium holds great significance as a solution paradigm for predicting the outcome of this game. A Nash equilibrium signifies a strategic profile wherein no individual player possesses the motivation to unilaterally deviate from their current strategy. It can be formally defined as Definition 1.
@@
If and only if a strategy profile $\mathcal{A}^{=}(a_{1}^{,}\dots,a_{M}^{*})$ satisfies the following condition, it can be considered as a Nash equilibrium of game $\Gamma$
@@
Next, we introduce an important framework called the exact potential game [50] through Definitions 2 and 3, to help us analyze the existence of Nash equilibrium and determine a Nash equilibrium solution for MUD-TOG.
@@
A game is regarded as an exact potential game if and only if a potential function $F(\mathcal{A}):\mathbb{A}\mapsto\mathbb{R}$ exists such that
$\small\begin{split}&U_{m}(a_{m},a_{-m})-U_{m}(a_{m}^{\prime},a_{-m})\\ &=F(a_{m},a_{-m})-F(a_{m}^{\prime},a_{-m}),\forall(a_{m},a_{-m}),(a_{m}^{\prime},a_{-m})\in\mathbb{A}.\end{split}$ (43)
@@
The exact potential game with finite strategy sets always has a Nash equilibrium and the finite improvement property (FIP) [50, 51].
The FIP implies that a Nash equilibrium can be obtained in a finite number of iterations by any best-response correspondence. Specifically, the best-response correspondence can be formally defined as follows:
@@
For each player $m\in\mathcal{M}$, their best response correspondence corresponds to a set-valued mapping $\mathbf{B}_{m}(a_{-m})$: $\mathbf{A}_{-m}\longmapsto\mathbf{A}_{m}$ such that
$\small\mathbf{B}_{m}(a_{-m})=\left\{a_{m}^{*}\mid a_{m}^{*}\in\underset{a_{m}\in\mathbf{A}_{m}}{\arg\max}U_{m}\left(a_{m},a_{-m}\right)\right\}.$ (44)
@@
The MUD-TOG is an exact potential game where the potential function $F(\mathcal{A})$ can be given as
@@
We can obtain the edge computing utility of UD $m$ by substituting (34) into (37) and (38), i.e.,
$U_{m}^{\mathrm{SUAV}}(\mathcal{A})=\frac{Q_{n}^{\mathrm{c}}}{V}E_{n,m}^{\mathrm{c}}+\sqrt{\frac{\gamma_{m}\eta_{m}D_{m}}{F_{n}^{\mathrm{max}}}}\sum_{i\in\mathbf{M}_{n}(\mathcal{A})}\sqrt{\frac{\gamma_{i}\eta_{i}D_{i}}{F_{n}^{\mathrm{max}}}}+$ (46)
$\sqrt{\frac{\gamma_{m}D_{m}+(1-\gamma_{m})p_{m}D_{m}}{B_{n}r_{n,m}}}\sum_{i\in\mathbf{M}_{n}(\mathcal{A})}\sqrt{\frac{\gamma_{i}D_{i}+(1-\gamma_{i})p_{i}D_{i}}{B_{n}r_{n,i}}}$
and
$U_{m}^{\mathrm{LUAV}}(\mathcal{A})=\sqrt{\frac{\gamma_{m}\eta_{m}D_{m}}{F_{u}^{\mathrm{max}}}}\sum_{i\in\mathbf{M}_{u}(\mathcal{A})}\sqrt{\frac{\gamma_{i}\eta_{i}D_{i}}{F_{u}^{\mathrm{max}}}}+$ (47)
$\sqrt{\frac{\gamma_{m}D_{m}+(1-\gamma_{m})p_{m}D_{m}}{B_{u}r_{u,m}}}\sum_{i\in\mathbf{M}_{u}(\mathcal{A})}\sqrt{\frac{\gamma_{i}D_{i}+(1-\gamma_{i})p_{i}D_{i}}{B_{u}r_{u,i}}}.$
Let $\beta_{s,m}=\sqrt{\frac{\gamma_{m}\eta_{m}D_{m}}{F_{s}^{\mathrm{max}}}}$ and $\phi_{s,m}=\sqrt{\frac{\gamma_{m}D_{m}+(1-\gamma_{m})p_{m}D_{m}}{Br_{s,m}}}$, where $m\in\mathcal{M}$ and $s\in\mathcal{S}$. Furthermore, for an arbitrary ordering of UDs, let us introduce the following notation:
$\small\begin{array}[]{ll}\beta_{s}^{\leq m}(\mathcal{A})=\underset{{\left\{j\in\mathbf{M}_{s}(\mathcal{A})\mid j\leq m\right\}}}{\sum}\beta_{j,s},&\beta_{s}^{>m}(\mathcal{A})=\underset{{\left\{j\in\mathbf{M}_{s}(\mathcal{A})\mid j>m\right\}}}{\sum}w_{j,s},\\ \phi_{s}^{\leq m}(\mathcal{A})=\underset{{\left\{j\in\mathbf{M}_{s}(\mathcal{A})\mid j\leq m\right\}}}{\sum}\phi_{j,s},&\phi_{s}^{>m}(\mathcal{A})=\underset{{\left\{j\in\mathbf{M}_{s}(\mathcal{A})\mid j>m\right\}}}{\sum}\phi_{j,s},\end{array}$
where $\mathbf{M}_{s}(\mathcal{A})$ represents the set of UDs who offload tasks to server $s$.
Suppose UD $m$ updates its current decision $a_{m}$ to the decision $a_{m}^{\prime}$ that leads to a change in its utility function, i.e., $U_{m}(a_{m},a_{-m})-U_{m}(a_{m}^{\prime},a_{-m})$. Based on the definition of the potential function, i.e., Definition 2, we demonstrate through the following four cases which also leads to an equal change in the potential function.
Case 1: Suppose that $a_{m}=0$ and $a_{m}^{\prime}=u$. According to (45), we can obtain the following conclusion
$F$ $(a_{m},a_{-m})-F(a_{m}^{\prime},a_{-m})$ (48)
$=$ $\ U_{m}^{\mathrm{loc}}-\beta_{u,m}\beta_{u}^{\leq m}(a_{m}^{\prime},a_{-m})-\phi_{u,m}\phi_{u}^{\leq m}(a_{m}^{\prime},a_{-m})$
$\ -\beta_{u,m}\beta_{u}^{>m}(a_{m}^{\prime},a_{-m})-\phi_{u,m}\phi_{u}^{>m}(a_{m}^{\prime},a_{-m})$
$=$ $\ U_{m}^{\mathrm{loc}}-U_{m}^{\mathrm{LUAV}}(a_{m}^{\prime},a_{-m})$
$=$ $\ U_{m}(a_{m},a_{-m})-U_{m}(a_{m}^{\prime},a_{-m}).$
Case 2: Suppose that $a_{m}=0$ and $a_{m}^{\prime}=n$, $\forall n\in\mathcal{N}$. According to (45), we can obtain the following conclusion
$F$ $(a_{m},a_{-m})-F(a_{m}^{\prime},a_{-m})$ (49)
$=$ $\ U_{m}^{\mathrm{loc}}-\beta_{u,m}\beta_{u}^{\leq m}(a_{m}^{\prime},a_{-m})-\phi_{u,m}\phi_{u}^{\leq m}(a_{m}^{\prime},a_{-m})$
$\ -\beta_{u,m}\beta_{u}^{>m}(a_{m}^{\prime},a_{-m})-\phi_{u,m}\phi_{u}^{>m}(a_{m}^{\prime},a_{-m})-\frac{Q_{n}^{\mathrm{c}}}{V}E_{n,m}^{\mathrm{c}}$
$=$ $\ U_{m}^{\mathrm{loc}}-U_{m}^{\mathrm{SUAV}}(a_{m}^{\prime},a_{-m})$
$=$ $\ U_{m}(a_{m},a_{-m})-U_{m}(a_{m}^{\prime},a_{-m}).$
Case 3: Suppose that $a_{m}=u$ and $a_{m}^{\prime}=n$, $\forall n\in\mathcal{N}$. According to (45), we can obtain the following conclusion
$F$ $(a_{m},a_{-m})-F(a_{m}^{\prime},a_{-m})$ (50)
$=$ $\ \beta_{u,m}\beta_{u}^{\leq m}(a_{m},a_{-m})+\phi_{u,m}\phi_{u}^{\leq m}(a_{m},a_{-m})$
$\ +\beta_{u,m}\beta_{u}^{>m}(a_{m},a_{-m})+\phi_{u,m}\phi_{u}^{>m}(a_{m},a_{-m})$
$-\beta_{n,m}\beta_{n}^{\leq m}(a_{m}^{\prime},a_{-m})-\phi_{n,m}\phi_{n}^{\leq m}(a_{m}^{\prime},a_{-m})$
$\ -\beta_{n,m}\beta_{n}^{>m}(a_{m}^{\prime},a_{-m})-\phi_{n,m}\phi_{n}^{>m}(a_{m}^{\prime},a_{-m})-\frac{Q_{n}^{\mathrm{c}}}{V}E_{n,m}^{\mathrm{c}}$
$=$ $\ U_{m}^{\mathrm{LUAV}}(a_{m},a_{-m})-U_{m}^{\mathrm{SUAV}}(a_{m}^{\prime},a_{-m})$
$=$ $\ U_{m}(a_{m},a_{-m})-U_{m}(a_{m}^{\prime},a_{-m}).$
Case 4: Suppose that $a_{m}=n$ and $a_{m}^{\prime}=n^{\prime}$, $\forall n,n^{\prime}\in\mathcal{N}$ and $n\neq n^{\prime}$. According to (45), we can obtain the following conclusion
$F$ $(a_{m},a_{-m})-F(a_{m}^{\prime},a_{-m})$ (51)
$=$ $\ \frac{Q_{n}^{\mathrm{c}}}{V}E_{n,m}^{\mathrm{c}}+\beta_{n,m}\beta_{n}^{\leq m}(a_{m},a_{-m})+\phi_{n,m}\phi_{n}^{\leq m}(a_{m},a_{-m})$
$\ +\beta_{n,m}\beta_{n}^{>m}(a_{m},a_{-m})+\phi_{n,m}\phi_{n}^{>m}(a_{m},a_{-m})$
$-\beta_{n^{\prime},m}\beta_{n^{\prime}}^{\leq m}(a_{m}^{\prime},a_{-m})-\phi_{n^{\prime},m}\phi_{n^{\prime}}^{\leq m}(a_{m}^{\prime},a_{-m})$
$\ -\beta_{n^{\prime},m}\beta_{n^{\prime}}^{>m}(a_{m}^{\prime},a_{-m})-\phi_{n^{\prime},m}\phi_{n^{\prime}}^{>m}(a_{m}^{\prime},a_{-m})-\frac{Q_{n^{\prime}}^{\mathrm{c}}}{V}E_{n^{\prime},m}^{\mathrm{c}}$
$=$ $\ U_{m}^{\mathrm{SUAV}}(a_{m},a_{-m})-U_{m}^{\mathrm{SUAV}}(a_{m}^{\prime},a_{-m})$
$=$ $\ U_{m}(a_{m},a_{-m})-U_{m}(a_{m}^{\prime},a_{-m}).$
Therefore, we can conclude that the MUD-TOG is an exact potential game.
@@
$\Gamma^{\prime}$ is also an exact potential game and has the same potential function as $\Gamma$.
@@
Since $\Gamma$ is an exact potential game, the equality (43) holds. Therefore, it remains valid if we restrict $(a_{m},a_{-m})$ and $(b_{m},a_{-m})$ to the new strategy space $\mathbb{A}^{\prime}$, which is a subset of $\mathbb{A}$. This provides evidence for the validity of this theorem.
The key idea of the MUD-TOG is to utilize the FIP to update the offloading strategies of the players iteratively until a Nash equilibrium is attained, which is outlined in Algorithm 1. The main steps are described as follows. i) All UDs choose local computing for the initial setting (Line 1). ii) Each iteration is partitioned into $N$ decision slots (Lines 4-15). Within each decision slot, one UD is chosen to update its offloading decision using the best response correspondence, while the offloading decisions of the remaining UDs remain unchanged. iii) If a lower utility is achieved and constraint (20c) is satisfied, the offloading decision of the UD is updated. Otherwise, the original offloading decision is retained. iv) When no UD changes its offloading decision, the MUD-TOG reaches the Nash equilibrium.

### v2

@@ V-A Stage 1: Task Offloading and Resource Allocation
Problem P1 is still a MINLP problem, and task offloading and resource allocation are coupled with each other. Considering that the UAVs are service providers in the considered UAV-assisted MEC system, we prioritize resource allocation strategies for the UAVs. Then, based on the resource allocation strategies, we optimize the offloading decisions of UDs.
V-A1 Resource Allocation
For any arbitrary task offloading decision profile $\mathcal{A}$ of the UDs, the servers decides resource allocation strategies to minimize problem P1. Define $z_{s,m}=\frac{F_{s,m}}{F_{s}^{\text{max}}}$ and remove irrelevant constant terms, the resource allocation subproblem can be formulated as
@@ V-A1 Resource Allocation
To solve problem P1.1, we first prove that the problem is a convex optimization problem through Lemma 1. Then, we can obtain the optimal resource allocation by Theorem 2.
@@
Refer to Section II of the supplementary material.
@@
Refer to Section III of the supplementary material.
V-A2 Task Offloading
@@
Therefore, we can design the utility function of UD $m$ as
@@
However, as shown in Eq. (30), the task offloading decision for UD $m$ is influenced by not only its individual demand, but also the offloading decisions made by other UDs. Considering the inherent competitive dynamics of task offloading among UDs, we utilize game theory to address this subproblem.
@@
  • $\mathbb{A}=\mathbf{A}_{1}\times\dots\times\mathbf{A}_{M}$ denotes the strategy space, wherein $\mathbf{A}_{m}=\{0,u\}\cup{\mathcal{N}}$ is the set of offloading strategies for player $m\ (m\in\mathcal{M})$, $a_{m}\in\mathbf{A}_{m}$ denotes the offloading decision of player $m$, and $\mathcal{A}=(a_{1},\dots,a_{M})\in\mathbb{A}$ denotes a strategy profile.
  • $(U_{m}(\mathcal{A}))_{m\in\mathcal{M}}$ is the utility function of player $m$ that assigns a real number to each strategy profile $\mathcal{A}$.
Every player strives to minimize its utility by selecting an appropriate strategy. Mathematically, the MUD-TOG can be described as the following distributed optimization problem
@@
(2) The Solution of MUD-TOG. We begin by introducing the concept of Nash equilibrium, which is commonly used to describe the solutions of MUD-TOG. A Nash equilibrium signifies a strategic profile wherein no individual player possesses the motivation to unilaterally change their current strategy, which can be formally defined as Definition 1.
@@
If and only if a strategy profile $\mathcal{A}^{*}=(a_{1}^{*},\dots,a_{M}^{*})$ satisfies the following condition, it can be considered as a Nash equilibrium of game $\Gamma$
@@
Next, we introduce an important framework called the exact potential game [49] through Definitions 23, and 4, to analyze whether there is a Nash equilibrium for MUD-TOG and how to obtain a Nash equilibrium.
@@
If the game $\Gamma$ has a potential function $F(\mathcal{A}):\mathbb{A}\mapsto\mathbb{R}$ that satisfies the following condition, it can be regarded as an exact potential game.
$\small\begin{split}&U_{m}(a_{m},a_{-m})-U_{m}(a_{m}^{\prime},a_{-m})=F(a_{m},a_{-m})-F(a_{m}^{\prime},a_{-m}),\\ &\forall(a_{m},a_{-m}),(a_{m}^{\prime},a_{-m})\in\mathbb{A}.\end{split}$ (38)
@@
A Nash equilibrium and the finite improvement path (FIP) property always exist for an exact potential game with finite strategy sets  [49, 50].
@@
A sequence of strategy profiles $\rho=(\mathcal{A}^{0},\mathcal{A}^{1},\mathcal{A}^{2},\ldots)$ is called an improvement path if for every index $i\geq 0$, $\mathcal{A}^{i+1}$ is obtained from $\mathcal{A}^{i}$ by allowing a player $m(i)$ (the single deviator in step $i$) to change their strategy and improve their utility, i.e., $U_{m(i)}(\mathcal{A}^{i+1})<U_{m(i)}(\mathcal{A}^{i})$. Furthermore, a game is said to have the FIP property if every improvement path in the game is of finite length.
The FIP property implies that any best-response correspondence can reach a Nash equilibrium in a finite number of iterations [49]. Definition 39 provides the formal definition of the best-response correspondence.
Definition 5.
For each player $m\in\mathcal{M}$, their best-response correspondence corresponds to a set-valued mapping $\mathbf{B}_{m}(a_{-m})$: $\mathbf{A}_{-m}\longmapsto\mathbf{A}_{m}$ such that
$\small\mathbf{B}_{m}(a_{-m})=\big{\{}a_{m}^{*}\mid a_{m}^{*}\in\underset{a_{m}\in\mathbf{A}_{m}}{\arg\min}U_{m}\left(a_{m},a_{-m}\right)\big{\}}.$ (39)
@@
The MUD-TOG is an exact potential game with the following potential function
@@
Refer to Section IV of the supplementary material.
@@
$\Gamma^{\prime}$ possesses the same potential function as $\Gamma$, which also is an exact potential game.
@@
Refer to Section V of the supplementary material.
The main idea for solving the MUD-TOG is to iteratively update the offloading strategies of the players using the best-response correspondence until a Nash equilibrium is reached. This process is outlined in Algorithm 1, and the main steps are described as follows. i) All UDs choose local computing for the initial setting (Line 1). ii) During each iteration, the process is partitioned into $N$ decision slots (Lines 4-15). In each slot, one UD is chosen to update its offloading decision based on the best-response correspondence, while the offloading decisions of the remaining UDs remain unchanged. iii) If a lower utility is achieved and constraint (21c) is satisfied, the offloading decision of the UD is updated. Otherwise, the offloading decision remains unchanged. iv) The MUD-TOG achieves a Nash equilibrium when no UD modifies its offloading decision.